a ball is thrown from the top of 10m high building with horizontal speed of 5m/s. (a) calculate the time it takes to drop to the floor, (b) its horizontal displacement at the point when it reaches the floor


1
Expert's answer
2021-02-03T16:11:57-0500

a)

"y=\dfrac{1}{2}gt^2,""t=\sqrt{\dfrac{2y}{g}}=\sqrt{\dfrac{2\cdot10\ m}{9.8\ \dfrac{m}{s^2}}}=1.43\ s."

b)

"x=v_0t=5\ \dfrac{m}{s}\cdot1.43\ s=7.15\ m."